OpenStudy (anonymous):

What is the projection of (2,6) onto (-1,5)? A.0.7(2,6) B.1.08(2,6) C.0.7(-1,5) D.1.08(-1,5)

OpenStudy (anonymous):

can't wait to see this ...

OpenStudy (anonymous):

Let\[x=\left[\begin{array} {c} 2 \\ 6 \end{array}\right]\]Then the projection of (2,6) onto (-1,5) is:\[\mbox{proj}(x)=(x\cdot u)u\]where the dot is the dot product (inner product), and u is a unit vector in the direction of (-1,5). So first you need to calculate the unit vector in the direction of (-1,5), then compute the dot product of x and u, and finally multiply that scalar times u itself.

OpenStudy (anonymous):

true :P

OpenStudy (anonymous):

lol :)

OpenStudy (anonymous):

\[Proj_b a = \frac{ a.b }{ |b| }.u_b = \frac{a.b}{|b|}.\frac{b}{|b|}\] where \[u_b\] is unit vector in direction of b

OpenStudy (anonymous):

Its D